# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ Appendix B Restricted Isometry Results
$\left|\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{\top}\right.$ $\left.\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathbb{E}\left[\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]\right|$
$=\left|\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathbb{E}\left[\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]\right.$
$\left.\quad+2\left(\mathbb{E}\left[\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]-\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right)^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})-2\left(\mathbb{E}\left[\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]-\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right)^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathbb{E}\left[\mathcal{S}_{\Omega}(\hat{\bm{Y}})\right]\right|$
@@
Bounding $B_{1}$: We will compute $B_{1}$ using the Hanson-Wright inequality, seen in Theorem A.3. We will first define $\bm{\xi}\in\mathbb{R}^{L}$ to be a Bernoulli random vector, where each entry is an i.i.d. Bernoulli random variable with parameter $p$, i.e. $\bm{\xi}_{\bm{\alpha}}=\xi_{\bm{\alpha}}$. Next, we define the following matrix $\bm{A}_{\hat{\bm{Y}}}\in\mathbb{R}^{\bm{L}\times L}$ as $(\bm{A}_{\hat{\bm{Y}}})_{\bm{\alpha}\bm{\beta}}=\langle\bm{Y},\hat{\bm{w}}_{\bm{\alpha}}\rangle\langle\bm{Y},\hat{\bm{w}}_{\bm{\beta}}\rangle$. We first remark that
$\bm{\xi}^{T}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})\bm{\xi}$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\xi_{\bm{\alpha}}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}\xi_{\bm{\beta}}$
@@
$=\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{T}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}}),$
@@
$(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\left(\xi_{\bm{\alpha}}-p\right)\left(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right)(\xi_{\bm{\beta}}-p)$
@@
$\mathcal{S}_{\Omega}^{C}(\hat{\bm{Y}})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}^{C}(\hat{\bm{Y}})$ $=(\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathbb{E}(\mathcal{S}_{\Omega}(\hat{\bm{Y}}))^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}(\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathbb{E}[\mathcal{S}_{\Omega}(\hat{\bm{Y}})])$
$=\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})-\mathcal{S}_{\Omega}(\hat{\bm{Y}})^{T}\bm{H}^{-1}_{\mathrm{offdiag}}\mathbb{E}[\mathcal{S}_{\Omega}(\hat{\bm{Y}})]$
$\quad-\mathbb{E}[\mathcal{S}_{\Omega}(\hat{\bm{Y}})]^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\hat{\bm{Y}})+\mathbb{E}[\mathcal{S}_{\Omega}(\hat{\bm{Y}})]^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathbb{E}[\mathcal{S}_{\Omega}(\hat{\bm{Y}})]$
@@
$=(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}]).$
As such, we will now proceed to use Theorem A.3 using $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$. We note that from Lemma B.4, setting $\bm{A}=\bm{B}=\frac{1}{\sqrt{n}}\bm{I}$ in the lemma statement, that $\|\bm{\xi}\|_{\psi_{2}}\leq C\sqrt{p}$ for some absolute constant $C>0$. Next, we compute that
@@
where the first inequality follows from the largest off-diagonal element of $\bm{H}^{-1}$ from Lemma A.8, the second inequality follows from adding a positive term to the sum, and the third inequality follows from Lemma A.5. Next, we will use a Gershgorin estimate to compute $\left\|\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right\|$ as follows:
@@
where the first inequality follows from Cauchy-Schwarz and the second inequality follows from Lemma A.8. Now, as $A_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}$ is diagonal-free and $\bm{\xi}-\mathbb{E}[\bm{\xi}]$ is a centered random vector, we can now say that for $t=Cp\max_{\bm{\alpha}}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\beta\log{n}$ for some sufficiently large constant $C>0$,
$\mathbb{P}\bigg{(}\bigg{|}(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $\bigg{|}>Cp\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\beta\log{n}\bigg{)}$
@@
as for both $\hat{\bm{w}}_{\bm{\alpha}}=\mathcal{P}_{\mathbb{T}}\bm{w}_{\bm{\alpha}}$ or $\hat{\bm{w}}_{\bm{\alpha}}=\bm{w}_{\bm{\alpha}}$ the minimum is achieved by the term on the right, from Lemma A.8.
Bounding $B_{2}$: The next step of this result requires bounding $B_{2}$. We will do this using the scalar Bernstein inequality, provided in Theorem A.2. To use this theorem, we need to decompose $B_{2}$ as a sum of independent random variables. To do this, notice that
@@
where the second inequality follows from Lemma A.8, and the final inequality is a numerical inequality.
@@
$\leq 4\left(\max_{\bm{\beta}}\|\hat{\bm{w}}_{\bm{\beta}}\|_{\mathrm{F}}^{2}\right)p\|\bm{Y}\|_{F}^{2}\lambda_{\max}(\hat{\bm{H}}),$
where the second inequality follows from Lemma A.5. Letting $t=\frac{p}{2}\sqrt{\frac{128\left(\max\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\right)\lambda_{\max}(\tilde{\bm{H}})\beta\log{n}}{3p}}$ for $\beta>1$, it follows from the scalar Bernstein inequality that, using the specified restriction $p\geq\frac{32}{3}\max_{\alpha}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\frac{\beta\log{n}}{\lambda_{\max}(\hat{\bm{H}})}$,
@@
We are now ready to prove Theorem 5.3.
B.1 Proof of Theorem 5.3
@@
The result follows from Lemmas B.6 and B.8, and the fact that $\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\leq\frac{1}{2}$ from Lemma A.8. ∎
@@
This second term can be analyzed in the same way as in the proof of Theorem 5.3, seen in Section B.1:
@@
Using the fact that $\|\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}=4$ and that $\lambda_{\max}(\bm{H})=2n$ from Lemma A.8, the result follows. ∎
@@
Let $\bm{Y},\bm{Z}\in\mathbb{R}^{n\times n}$ be any matrix with $\|\bm{Y}\|_{\mathrm{F}}=\|\bm{Z}\|_{\mathrm{F}}=1$. For a $r$, $\nu$-incoherent ground truth matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, and for $\mathcal{S}_{\Omega}$ defined as in (32) and $\bm{H}^{-1}_{\mathrm{offdiag}}$ defined as in (33), we have that for any $\beta>1$ and some absolute numerical constant $C>0$, if $p\geq\frac{8}{3}\beta\frac{\log{n}}{n}$, that
@@
This proof is similar to that of Lemma B.8, with some minor differences due to the asymmetry. Defining $\mathcal{S}_{\Omega}^{C}=\mathcal{S}_{\Omega}-\mathbb{E}[\mathcal{S}_{\Omega}]$, we have for any $\bm{Y},\bm{Z}$ that
@@
Bounding $T_{1}$: We will first bound $T_{1}$ using the Hanson-Wright inequality (Theorem A.3). First, we define the following matrix $\bm{A}_{\bm{Y},\bm{Z}}\in\mathbb{R}^{L\times L}$ as $(A_{\bm{Y},\bm{Z}})_{\bm{\alpha}\bm{\beta}}=\langle\bm{Z},\bm{w}_{\bm{\alpha}}\rangle\langle\mathcal{P}_{\mathbb{T}}\bm{Y},\bm{w}_{\bm{\beta}}\rangle$. As such, we can write, for fixed $\bm{Y},\bm{Z}$, the following:
$\bm{\xi}^{T}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})\bm{\xi}$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\xi_{\bm{\alpha}}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}\xi_{\bm{\beta}}$
@@
$=\mathcal{S}_{\Omega}(\bm{Z})^{T}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\mathcal{P}_{\mathbb{T}}\bm{Y}),$
@@
$(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\left(\xi_{\bm{\alpha}}-p\right)\left(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right)(\xi_{\bm{\beta}}-p)$
@@
$=\mathcal{S}_{\Omega}(\bm{Z})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\mathcal{P}_{\mathbb{T}}\bm{Y})-\mathcal{S}_{\Omega}(\bm{Z})^{T}\bm{H}^{-1}_{\mathrm{offdiag}}\mathbb{E}[\mathcal{S}_{\Omega}\mathcal{P}_{\mathbb{T}}\bm{Y}]$
@@
$=(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}]).$
With this equality established, we can now proceed with using Theorem A.3 to bound $T_{1}$. Using Lemma B.4 and setting $\bm{A}=\bm{B}=\frac{1}{\sqrt{n}}\bm{I}$ in the Lemma statement, we have that $\|\bm{\xi}\|_{\psi_{2}}\leq C\sqrt{p}$ for some absolute constant $C>0$. Next, we need to bound the Frobenius norm of $\|\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\|_{\mathrm{F}}^{2}$. To do this, notice that, for $\|\bm{Z}\|_{\mathrm{F}}=\|\bm{Y}\|_{\mathrm{F}}=1$,
@@
where the first inequality follows from Lemma A.8, the third inequality follows from Lemma A.5, and the final inequality follows from Lemmas  A.8 and A.6. Next, to bound $\|\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\|$, we will use a Gershgorin estimate as follows:
@@
where the first inequality follows from Gershgorin’s circle theorem, the second inequality follows from Cauchy Schwarz and Assumption 5.1, and the final inequality comes from Lemma A.8. Furthermore, as $\mathrm{Trace}(\bm{H}^{-1}_{\mathrm{offdiag}})=0$, $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])=0$. Taking $\beta>1$, for some sufficiently large constant $C>0$ we have from Theorem A.3 that
$\mathbb{P}\bigg{(}\big{|}$ $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{T}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])\big{|}>Cp\sqrt{\nu r}\frac{\beta\log{n}}{\sqrt{n}}\bigg{)}$
@@
Bounding $T_{2}$: To bound $T_{2}$ and $T_{3}$, we will use the scalar Bernstein inequality seen in Theorem A.2. We will bound $T_{2}$ first. Defining
@@
where the first inequality follows from the triangle inequality, the second follows from Cauchy-Schwarz, the third follows from Assumption 5.1, and the final inequality follows from Lemma A.8. Next, notice that
@@
where the second inequality follows from Assumption 5.1, the third inequality follows from Lemma A.8, the fourth inequality follows from Lemma A.5, and the final line follows from Lemma A.8. As such, we have that for $p\geq\frac{4}{3}\beta\frac{\log{n}}{n}$ that
@@
where the first inequality follows from the triangle inequality, the second follows from Cauchy-Schwarz, the third follows from Assumption 5.1, and the final inequality follows from Lemma A.8. Next, notice that
@@
where the second inequality follows from Assumption 5.1, the third inequality follows from Lemma A.8, the fourth inequality follows from Lemmas A.5 and  A.7, and the final line follows from Lemma A.6. As such, we have that for $p\geq\frac{8}{3}\beta\frac{\log{n}}{n}$ that
@@
Let $\Omega\subset\mathbb{I}$ be sampled with uniform Bernoulli probability $p$, and let $\mathbb{T}$ be the tangent space on $\mathcal{N}_{r}$ for a $r$, $\nu$-incoherent ground truth matrix $\bm{X}$. If $p\geq\frac{8}{3}\frac{\beta\log{n}}{{n}}$, then with probability at least $1-2n^{1-\beta}-6n^{-\beta}$ we have that, for some absolute constant $c>0$,
@@
Similarly to the proof of Theorem 5.3, seen in Section B.1, and in the proof of Lemma B.9, we can decompose the difference between $\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}$ and $\mathbb{E}[\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}]$ as concentration of $\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}-\mathbb{E}[\mathcal{F}_{\Omega}]\mathcal{P}_{\mathbb{T}}$ and the off-diagonal quadratic form term. As such, we can see that
@@
For the final result, if $p\geq\frac{C^{\prime}\beta\log{n}}{n}$ for some sufficiently large constant $C^{\prime}>0$, the conditions of Lemma B.9 hold with high probability and the expression can be simplified to

### v2

@@ Appendix B Restricted Isometry Results
[…]
@@
Bounding $B_{1}$: We will compute $B_{1}$ using the Hanson-Wright inequality, seen in A.3. We will first define $\bm{\xi}\in\mathbb{R}^{L}$ to be a Bernoulli random vector, where each entry is an i.i.d. Bernoulli random variable with parameter $p$, i.e. $\bm{\xi}_{\bm{\alpha}}=\xi_{\bm{\alpha}}$. Next, we define the following matrix $\bm{A}_{\hat{\bm{Y}}}\in\mathbb{R}^{\bm{L}\times L}$ as $(\bm{A}_{\hat{\bm{Y}}})_{\bm{\alpha}\bm{\beta}}=\langle\bm{Y},\hat{\bm{w}}_{\bm{\alpha}}\rangle\langle\bm{Y},\hat{\bm{w}}_{\bm{\beta}}\rangle$. We first remark that
$\bm{\xi}^{\top}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})\bm{\xi}$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\xi_{\bm{\alpha}}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}\xi_{\bm{\beta}}$
@@
[…]
@@
$(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\left(\xi_{\bm{\alpha}}-p\right)\left(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right)(\xi_{\bm{\beta}}-p)$
@@
[…]
@@
$=(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}]).$
As such, we will now proceed to use A.3 using $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$. We note that from B.4, setting $\bm{A}=\bm{B}=\frac{1}{\sqrt{n}}\bm{I}$ in the lemma statement, that $\|\bm{\xi}\|_{\psi_{2}}\leq C\sqrt{p}$ for some absolute constant $C>0$. Next, we compute that
@@
where the first inequality follows from the largest off-diagonal element of $\bm{H}^{-1}$ from A.8, the second inequality follows from adding a positive term to the sum, and the third inequality follows from A.5. Next, we will use a Gershgorin estimate to compute $\left\|\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right\|$ as follows:
@@
where the first inequality follows from Cauchy-Schwarz and the second inequality follows from A.8. Now, as $A_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}$ is diagonal-free and $\bm{\xi}-\mathbb{E}[\bm{\xi}]$ is a centered random vector, we can now say that for $t=Cp\max_{\bm{\alpha}}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\beta\log{n}$ for some sufficiently large constant $C>0$,
$\mathbb{P}\bigg(\bigg|(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\hat{\bm{Y}}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $\bigg|>Cp\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\beta\log{n}\bigg)$
@@
as for both $\hat{\bm{w}}_{\bm{\alpha}}=\mathcal{P}_{\mathbb{T}}\bm{w}_{\bm{\alpha}}$ or $\hat{\bm{w}}_{\bm{\alpha}}=\bm{w}_{\bm{\alpha}}$ the minimum is achieved by the term on the right, from A.8.
Bounding $B_{2}$: The next step of this result requires bounding $B_{2}$. We will do this using the scalar Bernstein inequality, provided in A.2. To use this theorem, we need to decompose $B_{2}$ as a sum of independent random variables. To do this, notice that
@@
where the second inequality follows from A.8, and the final inequality is a numerical inequality.
@@
[…]
where the second inequality follows from A.5. Letting $t=\frac{p}{2}\sqrt{\frac{128\left(\max\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\right)\lambda_{\max}(\tilde{\bm{H}})\beta\log{n}}{3p}}$ for $\beta>1$, it follows from the scalar Bernstein inequality that, using the specified restriction $p\geq\frac{32}{3}\max_{\alpha}\|\hat{\bm{w}}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\frac{\beta\log{n}}{\lambda_{\max}(\hat{\bm{H}})}$,
@@
B.1 Proof of Lemma 5.3
We are now ready to prove 5.3.
@@
The result follows from Lemmas B.6 and B.8, and the fact that $\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\leq\frac{1}{2}$ from A.8. ∎
@@
This second term can be analyzed in the same way as in the proof of 5.3, seen in Section B.1:
@@
Using the fact that $\|\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}=4$ and that $\lambda_{\max}(\bm{H})=2n$ from A.8, the result follows. ∎
@@
Let $\bm{Y},\bm{Z}\in\mathbb{R}^{n\times n}$ be any matrix with $\|\bm{Y}\|_{\mathrm{F}}=\|\bm{Z}\|_{\mathrm{F}}=1$. For a rank-$r$, $\nu$-incoherent ground truth matrix $\bm{X}$ with tangent space $\mathbb{T}$ on $\mathcal{N}_{r}$, and for $\mathcal{S}_{\Omega}$ defined as in (33) and $\bm{H}^{-1}_{\mathrm{offdiag}}$ defined as in (34), we have that for any $\beta>1$ and some absolute numerical constant $C>0$, if $p\geq\frac{8}{3}\beta\frac{\log{n}}{n}$, that
@@
This proof is similar to that of B.8, with some minor differences due to the asymmetry. Defining $\mathcal{S}_{\Omega}^{C}=\mathcal{S}_{\Omega}-\mathbb{E}[\mathcal{S}_{\Omega}]$, we have for any $\bm{Y},\bm{Z}$ that
@@
Bounding $T_{1}$: We will first bound $T_{1}$ using the Hanson-Wright inequality (A.3). First, we define the following matrix $\bm{A}_{\bm{Y},\bm{Z}}\in\mathbb{R}^{L\times L}$ as $(A_{\bm{Y},\bm{Z}})_{\bm{\alpha}\bm{\beta}}=\langle\bm{Z},\bm{w}_{\bm{\alpha}}\rangle\langle\mathcal{P}_{\mathbb{T}}\bm{Y},\bm{w}_{\bm{\beta}}\rangle$. As such, we can write, for fixed $\bm{Y},\bm{Z}$, the following:
$\bm{\xi}^{\top}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})\bm{\xi}$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\xi_{\bm{\alpha}}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})_{\bm{\alpha}\bm{\beta}}\xi_{\bm{\beta}}$
@@
$=\mathcal{S}_{\Omega}(\bm{Z})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\mathcal{P}_{\mathbb{T}}\bm{Y}),$
@@
$(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])$ $=\sum_{\begin{subarray}{c}\bm{\alpha},\bm{\beta}\in\mathbb{I}\\ \bm{\alpha}\neq\bm{\beta}\end{subarray}}\left(\xi_{\bm{\alpha}}-p\right)\left(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\right)(\xi_{\bm{\beta}}-p)$
@@
$=\mathcal{S}_{\Omega}(\bm{Z})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathcal{S}_{\Omega}(\mathcal{P}_{\mathbb{T}}\bm{Y})-\mathcal{S}_{\Omega}(\bm{Z})^{\top}\bm{H}^{-1}_{\mathrm{offdiag}}\mathbb{E}[\mathcal{S}_{\Omega}\mathcal{P}_{\mathbb{T}}\bm{Y}]$
@@
$=(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}]).$
With this equality established, we can now proceed with using A.3 to bound $T_{1}$. Using B.4 and setting $\bm{A}=\bm{B}=\frac{1}{\sqrt{n}}\bm{I}$ in the Lemma statement, we have that $\|\bm{\xi}\|_{\psi_{2}}\leq C\sqrt{p}$ for some absolute constant $C>0$. Next, we need to bound the Frobenius norm of $\|\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\|_{\mathrm{F}}^{2}$. To do this, notice that, for $\|\bm{Z}\|_{\mathrm{F}}=\|\bm{Y}\|_{\mathrm{F}}=1$,
@@
where the first inequality follows from A.8, the third inequality follows from A.5, and the final inequality follows from Lemmas  A.8 and A.6. Next, to bound $\|\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}}\|$, we will use a Gershgorin estimate as follows:
@@
where the first inequality follows from Gershgorin circle theorem, the second inequality follows from Cauchy Schwarz and 5.1, and the final inequality comes from A.8. Furthermore, as $\mathrm{Trace}(\bm{H}^{-1}_{\mathrm{offdiag}})=0$, $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])=0$. Taking $\beta>1$, for some sufficiently large constant $C>0$ we have from A.3 that
$\mathbb{P}\bigg(\big|$ $(\bm{\xi}-\mathbb{E}[\bm{\xi}])^{\top}(\bm{A}_{\bm{Y},\bm{Z}}\circ\bm{H}^{-1}_{\mathrm{offdiag}})(\bm{\xi}-\mathbb{E}[\bm{\xi}])\big|>Cp\sqrt{\nu r}\frac{\beta\log{n}}{\sqrt{n}}\bigg)$
@@
Bounding $T_{2}$: To bound $T_{2}$ and $T_{3}$, we will use the scalar Bernstein inequality seen in A.2. We will bound $T_{2}$ first. Defining
@@
where the first inequality follows from the triangle inequality, the second follows from Cauchy-Schwarz, the third follows from 5.1, and the final inequality follows from A.8. Next, notice that
@@
where the second inequality follows from 5.1, the third inequality follows from A.8, the fourth inequality follows from A.5, and the final line follows from A.8. As such, we have that for $p\geq\frac{4}{3}\beta\frac{\log{n}}{n}$ that
@@
where the first inequality follows from the triangle inequality, the second follows from Cauchy-Schwarz, the third follows from 5.1, and the final inequality follows from A.8. Next, notice that
@@
where the second inequality follows from 5.1, the third inequality follows from A.8, the fourth inequality follows from Lemmas A.5 and A.7, and the final line follows from A.6. As such, we have that for $p\geq\frac{8}{3}\beta\frac{\log{n}}{n}$ that
@@
Let $\Omega\subset\mathbb{I}$ be sampled with uniform Bernoulli probability $p$, and let $\mathbb{T}$ be the tangent space on $\mathcal{N}_{r}$ for a rank-$r$, $\nu$-incoherent ground truth matrix $\bm{X}$. If $p\geq\frac{8}{3}\frac{\beta\log{n}}{{n}}$, then with probability at least $1-2n^{1-\beta}-6n^{-\beta}$ we have that, for some absolute constant $c>0$,
@@
Similarly to the proof of 5.3, seen in Section B.1, and in the proof of B.9, we can decompose the difference between $\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}$ and $\mathbb{E}[\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}]$ as concentration of $\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}-\mathbb{E}[\mathcal{F}_{\Omega}]\mathcal{P}_{\mathbb{T}}$ and the off-diagonal quadratic form term. As such, we can see that
@@
For the final result, if $p\geq\frac{C^{\prime}\beta\log{n}}{n}$ for some sufficiently large constant $C^{\prime}>0$, the conditions of B.9 hold with high probability and the expression can be simplified to